# Grounding Text-To-Image Diffusion Models For Controlled High-Quality Image Generation

### v1

@@ Abstract
Large-scale text-to-image (T2I) diffusion models have demonstrated an outstanding performance in synthesizing diverse high-quality visuals from natural language text captions. Multiple layout-to-image models have been developed to control the generation process by utilizing a broad array of layouts such as segmentation maps, edges, and human keypoints. In this work, we present ObjectDiffusion, a model that takes inspirations from the top cutting-edge image generative frameworks to seamlessly condition T2I models with new bounding boxes capabilities. Specifically, we make substantial modifications to the network architecture introduced in ContorlNet to integrate it with the condition processing and injection techniques proposed in GLIGEN. ObjectDiffusion is initialized with pretraining parameters to leverage the generation knowledge obtained from training on large-scale datasets. We fine-tune ObjectDiffusion on the COCO2017 training dataset and evaluate it on the COCO2017 validation dataset. Our model achieves an AP50 of 46.6, an AR of 44.5, and a FID of 19.8 outperforming the current SOTA model trained on open-source datasets in all of the three metrics. ObjectDiffusion demonstrates a distinctive capability in synthesizing diverse, high-quality, high-fidelity images that seamlessly conform to the semantic and spatial control layout. Evaluated in qualitative and quantitative tests, ObjectDiffusion exhibits remarkable grounding abilities on closed-set and open-set settings across a wide variety of contexts. The qualitative assessment verifies the ability of ObjectDiffusion to generate multiple objects of different sizes and locations.
@@ 1 Introduction
Image generation is one of the most advanced research subfields of the relentlessly evolving Generative Artificial Intelligence (GenAI) and Deep Learning (DL). Thanks to its transformative practical applications, Generative Adversarial Networks (GANs) [14] have succeeded in drawing considerable attention both in academia and industry. Diffusion-based powerful image generation models [1, 38, 50, 53] are capable of synthesizing stunningly detailed realistic images of high-quality that capture diverse visual concepts.
@@
One major challenge that hinders the widespread adoption of generative AI in general and image generation models in particular is the difficulty of customization. Customization is an active research topic that aims to steer the image generation process towards producing controllable content that adheres to user-specified requirements. While text prompts [32, 41] are the most common type of input, it is deemed inadequate in cases that necessitate a high degree of content and style control in the synthesized image. Particularly, the image caption falls short in providing a precise description of the sizes, colors, shapes, textures, spatial alignments, and interactions between the objects present in the synthesized image. This limitation prompted computer vision researchers to adopt and integrate more effective types of control formats, in conjunction with the naive caption to achieve a greater degree of controllability over the output scenes. Segmentation maps [76], normal maps [64], depth maps [48], edge maps [4, 70], human keypoints [5], bounding boxes [49], user scribbles, cartoon drawings, image styles or a combination of them are among the well-established conditioning modalities that can be leveraged to tailor the image generation.
@@
In this work, we present ObjectDiffusion, a conditional image generative framework that augments text-to-image models (T2I) [36, 47] by incorporating new control capabilities. ObjectDiffusion can condition T2I models on object detection [49] annotations. That is, in addition to the text prompts, we condition T2I models on object labels and their corresponding bounding boxes. The object labels are not limited to a predefined set of categories, rather they can be in open-ended text format, for example, “a stripped pink full face motorcycle helmet”. ObjectDiffusion consists of two pretrained networks. The first is a locked text-to-image Stable Diffusion (SD) [50] model pretrained on filtered high-resolution images from the LAION-5B [54] dataset. The second module is GroundNet, a parallel trainable network consisting of pretrained encoding blocks and a middle block of GLIGEN [28]. GroundNet is pretrained on Object365 [56], GoldG (Flickr and VG) [26], SBU [39], and CC3M [57] datasets but not fine-tuned on the COCO [30] dataset. GroundNet is integrated into the locked Stable Diffusion [50] base model via zero convolutional layers.
@@
Our approach integrates the conditional architecture outline of ControlNet [73] with the grounding method established in GLIGEN by seamlessly leveraging their design concepts and techniques to advance the grounding capabilities while increasing the image quality. Specifically, the architecture of our proposed framework is inspired from ControlNet to optimally leverage the pretraining knowledge. However, since ControlNet does not support our conditional layout, the architecture and initialization of GroundNet are adopted from GLIGEN to enable processing and integrating the information of the object names and bounding boxes. We fine-tune ObjectDiffusion on the COCO2017 [30] training dataset and conduct a quantitative and qualitative evaluation on the COCO2017 [30] validation dataset.
@@
  • We propose a cutting-edge framework and demonstrate its abilities in grounding pretrained text-to-image models on open-ended object descriptions and bounding box annotations.
  • Our model ObjectDiffusion outperforms existing state-of-the-art models trained on open-source datasets on the FID [17] score, AP50 score, and AR [44] score. Our framework is capable of generating controlled images with superior quality.
  • We reduce the training cost substantially by effectively leveraging the generation and grounding capabilities of two pretrained models, SD and GLIGEN.
  • We evaluate the performance of our model and analyze the results of the quantitative and qualitative evaluations.
@@ 2.1 Diffusion Models
In 2015, Sohl-Dickstein et al. [58] established the theoretical foundation of the probabilistic diffusion model, a framework that works by iteratively adding random noise to the ground-truth images until all the structure is lost, and then learns how to denoise the perturbed images to restore their original structures through recursive runs of noise prediction neural network $\epsilon_{\theta}({x}_{t},t)$ over $T$ time-steps. Alternative to noise prediction, score-based [59] generative models $\mathbf{s}_{\boldsymbol{\theta}}({x})$ sample new instances by estimating the gradients of the log probability distribution of the data $\nabla_{{x}}\log p_{\text{data }}({x})$. Diffusion models and score-based generative models can be formalized under a unified framework [60].
Dhariwal et al. [8] demonstrated that probabilistic diffusion models [58] have outperformed the prominent GANs [14] both in quality and diversity metrics. Rombach et al. [50] presented Stable Diffusion (SD), a remarkable work that denoises the image in latent space in order to reduce the computational and memory requirements of operating in pixel-level space.
@@ 2.2 Control Layout Overview
Conditional models [72, 28, 68, 73, 45, 75, 35] empower image generative models, for the most part text-to-image models, to produce controllable outputs by extending them to accept single or multiple external layout maps, generally called conditions. Scribbles and sketches are simple and vague types of conditions that can be created conveniently by users. Segmentation maps [76], on the other hand, offer a granular control over the image content by dictating a pixel-level condition map of the image lattice. Hough lines [15] and Canny edges [4] can condition generative models to capture subtle texture and fine-grained details that can not be specified via bounding boxes and segmentation maps. Keypoints [5] is a format that allows creating images from skeletal structures that likely depict human postures. Advanced types of control formats include normal maps [64] and depth maps [48] both of which challenge the generative models by widening the conditioning task scope to integrate information about the perceived depth of images.
@@ 2.3 Conditional Image Generative Models
T2I models utilizes image caption to guide the generation process. Although simple and powerful, text prompt alone is insufficient for tailoring the generation process into creating customized outputs. To address this limitation, various layout-to-image systems [21, 29, 61, 63, 71, 74] have been proposed to use more advanced conditional modalities to enable a fine-grained control. Conditional models render complex images with style and content complying with the control layout. PITI [67] is a layout-to-image diffusion model that is trained from scratch on conditional tokens. Training generative models from scratch is inefficient and entails a very high training cost. T2I-Adopters [35] are light-weight neural network blocks borrowed from NLP [19] for facilitating fine-tuning huge T2I models on various types of guiding inputs like depth, segmentation, keypoints, sketch, and color maps. Some diffusion-based methods like ReCo [72] fine-tune a pretrained text-to-image model on bounding box and open-ended text information, however this continuous learning may cause a catastrophic loss of generation knowledge [20, 52].
To circumvent this, GLIGEN [28] proposed to freeze the weights of a pretrained SD model [50] and insert trainable gated self-attention layers [65] in the attention blocks of the frozen SD to handle the grounding tokens. InstanceDiffusion [68] takes the grounding task one step further by proposing a framework for precise object-level control while simultaneously enhancing the image fidelity and the multi-object sampling capability. Zhang et al. [73] goes further by introducing ControlNet [73], a groundbreaking general-purpose conditional model. It achieves this by freezing the pretrained model weights and training a copy of the pretrained locked backbone model. The locked model is preceded by a sequence of convolutional layers for mapping the conditions into the trainable network. UniControl [45] can simultaneously condition on an array of conditions by embedding all task instructions using a task-aware HyperNet [16, 66]. UniControl demonstrates zero-shot conditioning capabilities of unseen control types. Uni-ControlNet builds on ControlNet to introduce unified frameworks to simultaneously handle multiple control formats. Uni-ControlNet [75] enables training and inferring with a composite of modalities without the need for separate training for each modality.
@@ 2.4 Training-Free Conditional Methods
Traditional conditional models require computationally expensive training on a dataset of image-control pairs. Training-free [11, 12, 27, 29, 61, 71, 74] methods circumvent this by performing the conditioning on pretrained T2I models during the sampling process eliminating the requirement for training or fine-tuning. BoxDiff [69] achieves the positional alignment effect by manipulating specific regions of the cross-attention map that correspond to the constraining boundary box regions. Guidance-based methods can achieve control by designing task-specific guidance procedures. For example, Layout Control Guidance [6] applies forward and backward guidance losses on the attention map to steer the sampling towards generating object-level control. Universal Guidance [3] offers a solution to avoid training a network on noisy data by leveraging available pretrained models like semantic segmentation and object recognition neural networks. The goal is to carefully design and minimize the loss function $\ell\left(c,f\left(\hat{z}_{0}\right)\right)$ between the guidance function $f(\cdot)$ of the latent image $z$ and the control module $c$.
Despite their success, training-free control approaches prolong sampling time and decrease perceptual quality, and as a result, they are usually used in combination with regular training-based conditional models.
@@ 3 Method
For controllable image generation, we introduce ObjectDiffusion, a framework that takes inspirations from two remarkable conditional models GLIGEN [28] and ControlNet [73] to seamlessly augment open-source text-to-image generation diffusion models on control layout. Concretely, we combine the powerful network architecture of ControlNet with the condition processing and injection techniques proposed in GLIGEN. Our conditional layout is the object detection annotations consisting of rectangular boxes that specify the locations of objects on the image and the associated open-ended object text that describe the content inside each of the bounding boxes [72]. ObjectDiffusion incorporates the grounding information described above with a single image caption to generate high-fidelity images where objects specified by detection labels are rendered in the regions outlined by the bounding boxes. In this section, we explain the control layout and break down the encoding techniques used to process and integrate the condition inputs into the baseline model. We represent the overall model architecture and articulate the learning objective.
Figure 1: The grounding tokens consist of the concatenation of the CLIP [46] encoded object entities and their Fourier [34] embedded bounding boxes.Entity names do not need to be present in the caption, for example, the “car” is not mentioned in the caption. The figure is adopted and modified from [28].
@@ 3.1 Entity Grounding Layout
Due to its simplicity and effectiveness in controlling the generation process, we focus on conditioning diffusion-based models on object detection annotations. The object detection serves as a condition to ground text-to-image frameworks into rendering specific objects in specified regions of the synthesized images. In particular, our control layout consists of a set of grounding tokens, each of which is composed of one-to-one pairs of semantic and spatial information. The semantic instruction is a textual description of the object, for example, “A shiny luxury red sports car”. Each object in the image can be defined in an open-ended text format [72], rather than being limited to a predefined set of labels like COCO categories [30]. For spatial instruction, each entity is attributed by a single 2D rectangular box which determines the location of the object in the image. The box is represented in the form of an object detection rectangular box, that is, a list of two points needed to construct the detection box $[x_{1},y_{1},x_{2},_{1}]$ where $(x_{1},y_{1})$ is the upper-left point and $(x_{2},y_{2})$ is the lower-right point. While the objects of the image are defined by the control layout, the overall image outline is specified by a single global image caption, similar to T2I models.
Figure 2: ObjectDiffusion architecture is divided into two parallel networks. First, we utilize Stable Diffusion 1.4v [50] as a pretrained text-to-image model depicted with blue color. The second is a trainable GroundNet which consists of the encoder and middle blocks from GLIGEN [28] represented with red color. GroundNet is capable of processing the conditional layout $g$ which consists of the encoded positional and semantic tokens. During training, both networks receive the time $t$, the caption $c$ and the image input $z_{t}$. The two networks are connected via zero-convolutional layers, highlighted with gray color. ObjectDiffusion operates in latent space. The Image Encoder and Image Decoder project the input image from pixel space to latent space, and the output image from latent space back to pixel space, respectively. Our model architecture design is inspired by ControlNet [73].
3.2 Fusing Conditional Entities Overview
Processing Grounding Semantic Tokens: Conditional semantic tokens are strings of varying length consisting of single or multiple words. These object-level text captions are processed similar to the regular image caption. Specifically, they are encoded using a pretrained CLIP ViT-L/14 [46] text encoder. The CLIP encoder denoted as $\mathcal{E}_{\text{text}}(e)$ embeds the entity $e$ into a tensor feature of length 768.
Processing Grounding Spatial Tokens: We embed the spatial information into an expressive periodic tensor via a Fourier embedding [34] block. The mathematical representation of Fourier transformation $F(\cdot)$ of a box $b=[x_{1},y_{1},x_{2},y_{2}]$ is expressed as the following:
@@ 3.2 Fusing Conditional Entities Overview
where $M$ is the number of frequencies, and $F_{i}(b)$ is Fourier transformation at the $f_{i}$ frequency. The output of embedding $F(b)$ is 4 (box coordinates) * 8 (frequency) * 2 (sin & cos) = 64-dimensional tensor per bounding box. The Fourier embedder does not define any trainable parameters.
Concatenating Conditional Features: For each of the grounding entity, we first concatenate the encoded text entity $\mathcal{E}_{\text{text}}(e)$ with the embedded positional information $F(b)$ into a single tensor of size 832. Then we feed this tensor as an input to a Multilayer Perceptron (MLP) [43] consisting of three fully connected linear layers of sizes (832, 515, 768) and a Sigmoid-weighted Linear Unit (SiLU) [10] :
@@
The MLP transforms the concatenated grounding feature vectors from 832 to a transformer vector $o_{i}$ of standard size 768. Assuming that the maximum number of entities allowed per image is a hyperparameter $N$, the final control vector $o$ has the size $(N,768)$. Figure 1 illustrates the processing of grounding information and caption.
Integrating Conditional Features: To inject the extracted control features $o$ into the standard transformer blocks of LDM [50], we follow GLIGEN and insert a new gated self-attention $\operatorname{SA}\left(\cdot\right)$ layer between the self-attention and the cross-attention of the transformer block [2]. The gated self-attention layer takes both the intermediate latent visual features $v$ and the external control features $o$ as inputs and integrates them as shown in the equation:

### v2

@@ Abstract
[…]
@@ 1 Introduction
Image generation is a prominent subfield within the rapidly evolving domains of Generative Artificial Intelligence (GenAI) and Deep Learning (DL). Due to their transformative practical applications, Generative Adversarial Networks (GANs) [14] have attracted considerable attention in both academia and industry. Diffusion-based image generation models [1, 38, 50, 54] have emerged as powerful base models, capable of synthesizing stunningly detailed realistic images of high-quality, spanning a diverse array of concepts and domains.
@@
One major challenge hindering the widespread adoption of generative AI, in general, and image generation models, in particular, is the difficulty of customization. Customization is an active research area focused on steering the image generation process towards producing controllable content that meets user-specified requirements. While text prompts [32, 41] are the most common type for controlling image generation models, they are often deemed inadequate when a high degree of control over content and style is required in the synthesized image. Specifically, text captions fall short in providing precise descriptions of the sizes, colors, shapes, textures, spatial alignments, and interactions between objects in the generated image. This limitation has prompted the adaptation and integration of more effective control formats, alongside the caption, to achieve a greater degree of controllability over the output scenes. Well-established conditioning modalities, such as segmentation maps [77], normal maps [65], depth maps [48], edge maps [4, 71], human keypoints [5], bounding boxes [49], user scribbles, cartoon drawings, image styles, or combinations thereof, have been leveraged to tailor image generation.
@@
In this work, we present ObjectDiffusion, a conditional image generative model that augments text-to-image (T2I) models [36, 47] by incorporating new control modalities. ObjectDiffusion can condition T2I models on object detection annotations [49]. Specifically, in addition to text prompts, we ground T2I models using object names and their corresponding bounding boxes. The object names are not limited to a predefined set of categories; rather, they can be in an open-ended text format, for example, a striped pink full-face motorcycle helmet. ObjectDiffusion consists of two pretrained networks. The first is a fixed T2I Stable Diffusion (SD) [50] model, pretrained on filtered high-resolution images from the LAION-5B [55] dataset. The second module is GroundNet, a parallel, trainable network that consists of the pretrained encoding blocks and the middle block of GLIGEN [28]. GroundNet is pretrained on the Object365 [57], GoldG (Flickr and VG) [26], SBU [39], and CC3M [58] datasets but is not fine-tuned on the COCO [30] dataset. GroundNet is integrated into the locked SD [50] base model via zero convolution layers.
@@
Our approach integrates the conditional architecture design of ControlNet with the grounding method established in GLIGEN by seamlessly leveraging their design concepts and techniques to enhance grounding capabilities while improving image quality. The overall architecture of our proposed model is inspired by ControlNet, allowing optimal use of pretrained knowledge while enabling fine-tuning with new control inputs. However, since ControlNet does not support the conditional layout of entity names and bounding boxes, the architecture and initialization of GroundNet are adopted from GLIGEN to process and integrate the object names and bounding boxes information. We fine-tune ObjectDiffusion on the COCO2017 [30] training dataset and conduct both quantitative and qualitative evaluations on the COCO2017 [30] validation dataset.
@@
  • We propose a cutting-edge model, ObjectDiffusion, and demonstrate its ability to ground pretrained T2I models using open-ended object descriptions and bounding box annotations.
  • ObjectDiffusion combines the network architecture of ControlNet [74] with the grounding technique of GLIGEN [28], enhancing both image quality and controllability.
[…]
  • We significantly reduce training costs by effectively leveraging the generation and grounding capabilities of two pretrained models: SD [50] and GLIGEN [28].
  • We evaluate the performance of our model through both quantitative and qualitative assessments, providing a comprehensive analysis of the results.
@@ 2.1 Diffusion Models
Sohl-Dickstein et al. [59] established the theoretical foundation of probabilistic diffusion models, a framework in which random noise is iteratively added to ground-truth images until all structure is lost. The model then learns to denoise the perturbed images and restore their original structure through recursive applications of the noise prediction neural network $\epsilon_{\theta}({x}t,t)$ over $T$ time-steps. In contrast to noise prediction, score-based generative models [60] $\mathbf{s}{\boldsymbol{\theta}}({x})$ sample new instances by estimating the gradients of the log probability distribution of the data, $\nabla_{{x}}\log p_{\text{data }}({x})$. Diffusion models and score-based generative models can be formalized under a unified framework [61].
Dhariwal et al. [8] demonstrated that probabilistic diffusion models [59] outperform prominent GANs [14] in both quality and diversity metrics. Rombach et al. [50] introduced Stable Diffusion (SD), a groundbreaking approach that denoises images in latent space to reduce the computational and memory requirements associated with operating in pixel-level space. Stable Diffusion converts images to latent space via an image encoder and returns them to pixel space through an image decoder.
@@ 2.2 Control Layout Overview
Conditional models [73, 28, 69, 74, 45, 76, 35] empower image generative models, particularly text-to-image models, to produce controllable outputs by extending them to accept a single or multiple external layout maps, commonly referred to as conditions. Scribbles and sketches are simple, vague types of conditions that can be easily created by users. In contrast, segmentation maps [77] offer granular control over the image content by specifying a pixel-level condition map of the image lattice. Hough lines [15] and Canny edges [4] condition generative models to capture subtle textures and fine-grained details that cannot be specified through bounding boxes and segmentation maps. Keypoints [5] provide a format for generating images based on skeletal structures, which likely depict human postures. Advanced control formats include normal maps [65] and depth maps [48], both of which broaden the scope of the conditioning task by incorporating information about the perceived depth of images.
@@ 2.3 Conditional Image Generative Models
T2I models utilize image captions to guide the generation process. Although simple and effective, text prompts alone are insufficient for tailoring the generation process to produce customized outputs. To address this limitation, various layout-to-image models [21, 29, 62, 64, 72, 75] have been proposed, which employ more advanced conditional modalities, outlined in Section 2.2, to enable fine-grained control. Conditional models generate complex images with style and content that comply with the provided control layout. PITI [68] is a layout-to-image diffusion model trained from scratch on conditional tokens. Training generative models from scratch is inefficient and incurs high training costs, as it does not benefit from the vast generation knowledge available in pretrained T2I models. To harness the pretraining knowledge, T2I-Adopters [35] are lightweight neural network blocks adapted from NLP [19] that facilitate the fine-tuning of large T2I models on various types of guiding inputs, such as depth, segmentation, keypoints, sketches, and color maps.
ReCo [73] fine-tunes a pretrained T2I model using bounding box and open-ended text information. However, this continuous learning process may result in a catastrophic loss of generation knowledge [20, 53]. To mitigate this issue, GLIGEN [28] proposes freezing the weights of a pretrained SD model [50] and inserting trainable gated self-attention layers [66] into the attention blocks of the frozen SD model to handle the external grounding input. InstanceDiffusion [69] advances the grounding task further by introducing a model for precise object-level control while simultaneously enhancing image fidelity and multi-object sampling capabilities.
Zhang et al. [74] goes further by introducing ControlNet [74], a groundbreaking general-purpose conditional model. They achieve this by freezing the weights of the pretrained model and training a copy of the pretrained locked backbone model. The locked model is preceded by a sequence of convolution layers that map the conditions into the trainable network. UniControl [45] can simultaneously condition on an array of conditions by embedding all task instructions using a task-aware HyperNet [16, 67]. UniControl demonstrates zero-shot conditioning capabilities for previously unseen control types. Uni-ControlNet builds on ControlNet by introducing a unified framework to simultaneously handle multiple control formats. Uni-ControlNet [76] enables training and inference with a composite of modalities without the need for separate training for each modality.
@@ 2.4 Training-Free Conditional Methods
Traditional conditional models require computationally expensive training on datasets of image-control pairs. Training-free methods [11, 12, 27, 29, 62, 72, 75] circumvent this by performing conditioning on pretrained T2I models during the sampling process, thereby eliminating the need for training or fine-tuning. BoxDiff [70] achieves positional alignment by manipulating specific regions of the cross-attention map that correspond to constraining bounding box regions. Guidance-based methods can achieve control by designing task-specific guidance procedures. For example, Layout Control Guidance [6] applies forward and backward guidance losses on the attention map to steer the sampling process towards generating object-level control. Universal Guidance [3] offers a solution to avoid training a network on noisy data by leveraging available pretrained models, such as semantic segmentation and object recognition neural networks. The goal is to carefully design and minimize the loss function $\ell\left(c,f\left(\hat{z}_{0}\right)\right)$ between the guidance function $f(\cdot)$ of the latent image $z$ and the control module $c$.
Despite their success, training-free control approaches can prolong sampling time and decrease perceptual quality. As a result, they are typically used in conjunction with regular training-based conditional models.
@@ 3 Method
We introduce ObjectDiffusion, a model designed for grounded image generation based on object names and their locations. ObjectDiffusion draws inspiration from two notable conditional models, GLIGEN [28] and ControlNet [74], to seamlessly condition text-to-image diffusion models using a control layout. Specifically, we modify and combine the powerful network architecture of ControlNet with the condition processing and injection techniques proposed in GLIGEN. Our grounding layout consists of object detection annotations, where each object is defined by an open-ended text description [73], rather than a label as in conventional object detection tasks, along with a bounding box that specifies the location of the object in the image. ObjectDiffusion incorporates the grounding information described above, along with a single image caption, to generate high-fidelity images where objects specified by the entity names are rendered in the regions outlined by the bounding boxes. In this section, we explain the control layout and break down the encoding techniques used to process and integrate the condition inputs into the baseline model. We also present the overall model architecture and articulate the learning objective.
Figure 1: The grounding tokens are formed by fusing the CLIP [46] encoded object entities with their Fourier [34] embedded bounding boxes. The concatenation vector is processed through an MLP [43], which produces a standard-sized output vector of 768.
@@ 3.1 Entity Grounding Layout
Due to its simplicity and effectiveness in controlling the generation process, we focus on grounding diffusion-based models on object detection annotations. The object detection serves as a condition to ground T2I models into rendering specific objects in specified regions of the synthesized images. In particular, our control layout consists of a set of grounding tokens, each of which is composed of one-to-one pairs of semantic and spatial information. The semantic instruction is a textual description of the object, for example, a shiny luxury red sports car. Each object in the image can be defined in an open-ended text format [73], rather than being limited to a predefined set of labels like COCO categories [30]. For spatial instruction, each entity is attributed by a single 2D rectangular box which determines the location of the object in the image. The box is represented in the form of an object detection rectangular box defined by the two points, the upper-left point $(x_{1},y_{1})$ and the lower-right point $(x_{2},y_{2})$, needed to construct the detection box $[x_{1},y_{1},x_{2},y_{2}]$. While the objects of the image are defined by the control layout, the overall image outline is specified by a single global image caption, similar to T2I models.
Figure 2: ObjectDiffusion architecture is divided into two parallel networks. For the first network, we utilize Stable Diffusion 1.4v [50] as a pretrained text-to-image model, depicted in blue. The second network is a trainable GroundNet, which consists of the encoder and middle blocks from GLIGEN [28], represented in red. GroundNet injects the encoded conditional layout $g$, which consists of the positional and semantic tokens. During training, both networks receive the time $t$, the caption $c$, and the latent noised image input $z_{t}$. The two networks are connected via zero-convolution layers, highlighted in gray. ObjectDiffusion operates in latent space. The Image Encoder and Image Decoder project the input image from pixel space to latent space, and the output image from latent space back to pixel space, respectively. Our model architecture design is inspired by ControlNet [74].
3.2 Grounding Method Overview
Processing Semantic Grounding Tokens: Conditional semantic tokens are strings of varying lengths consisting of one or more words. These object-level text captions are processed similarly to regular image captions. Specifically, they are encoded using a pretrained CLIP ViT-L/14 [46] text encoder. The CLIP encoder, denoted as $\mathcal{E}_{\text{text}}(e)$, embeds the entity $e$ into a tensor feature of length 768.
Processing Spatial Grounding Tokens: We embed the spatial information, presented as bounding boxes, into an expressive periodic tensor via a Fourier embedding [34] block. The mathematical representation of the Fourier transformation $F(\cdot)$ of a box $b=[x_{1},y_{1},x_{2},y_{2}]$ is expressed as follows:
@@ 3.2 Grounding Method Overview
where $M$ is the number of frequencies, and $F_{i}(b)$ is the Fourier transformation at the $f_{i}$ frequency. The output of the embedding $F(b)$ is 4 $\times$ 8 $\times$ 2 (box coordinates $\times$ frequency $\times$ sin & cos), which results in a 64-dimensional tensor per bounding box. The Fourier embedder does not add any trainable parameters.
Fusing Conditional Features: For each grounding entity, we first concatenate the encoded text entity $\mathcal{E}_{\text{text}}(e)$ with the embedded positional information $F(b)$ into a single tensor of size 832. Then, we feed this tensor as input to a Multilayer Perceptron (MLP) [43] consisting of three fully connected linear layers of sizes (832, 512, 768) and a Sigmoid-weighted Linear Unit (SiLU) [10]:
@@
The MLP transforms the concatenated grounding feature vectors from 832 to a transformer vector $o_{i}$ of standard size 768. Assuming that the maximum number of entities allowed per image is a hyperparameter $N$, the final control vector $o$ has size ($N$, 768). Figure 1 illustrates the processing and fusion of the grounding inputs.
Integrating Conditional Features: To inject the extracted control features $o$ into the standard transformer blocks of LDM [50], we follow GLIGEN and insert a new gated self-attention $\operatorname{SA}\left(\cdot\right)$ layer between the self-attention and the cross-attention of the transformer block [2]. The gated self-attention layer takes both the intermediate latent visual features $v$ and the external control features $o$ as inputs and integrates them, as shown in the equation: